# Solving a Stackelberg Game on Transportation Networks in a Dynamic Crime Scenario: A Mixed Approach on Multi-Layer Networks

### v1

@@ Abstract
Interdicting a criminal with limited police resources is a challenging task as the criminal changes location over time. The size of the large transportation network further adds to the difficulty of this scenario. To tackle this issue, we consider the concept of a layered graph. At each time stamp, we create a copy of the entire transportation network to track the possible movements of both players, the attacker and the defenders. We consider a Stackelberg game in a dynamic crime scenario where the attacker changes location over time while the defenders attempt to interdict the attacker on his escape route. Given a set of defender strategies, the optimal attacker strategy is determined by applying Dijkstra’s algorithm on the layered networks. Here, the attacker aims to minimize while the defenders aim to maximize the probability of interdiction. We develop an approximation algorithm on the layered networks to find near-optimal strategy for defenders. The efficacy of the developed approach is compared with the adopted MILP approach. We compare the results in terms of computational time and solution quality. The quality of the results demonstrates the need for the developed approach, as it effectively solves the complex problem within a short amount of time.
@@
Resource allocation, Multi-layer time expanded network, Vehicle routing and scheduling, Network optimization, Exact algorithm, Approximation algorithm
@@ 1 Introduction
We consider a Stackelberg game where the defender is the leader, and the attacker is the follower. The assumption is that the defenders have information only about the crime location in a large transportation network. In this Stackelberg game, first, the defenders commit to a strategy, and then the attacker generates the best response to the given defender strategy. We use the concept of a multi-layer network in which, at each time-stamp, we create a copy of the entire network. We consider edge length as the time factor to create connections between these multi-layer networks. This means that depending on the edge length, we choose the layers from which the start and end nodes of that particular edge are selected. In this way, for all edges in the original transportation network, we create corresponding edges in the multi-layer network. We consider a set of strategies for both players, i.e., the defenders and the attacker. For the defender, we use a mixed strategy set where each strategy is assigned a mixed probability, with the sum of these probabilities equaling one. For the attacker, we consider a pure strategy set.
@@
To generate the optimal attacker strategy, we first assign weights to the nodes depending on the mixed probabilities and the nodes’ presence in the defender strategies. Considering all defender strategies, we sum up the probabilities if any node is present in multiple defender strategies. Then, we assign the weight of each node to all of its incoming edges. Finally, we apply Dijkstra’s algorithm on this network to find the optimal strategy for the attacker. For the defender, we create a multi-layer network in the same way as for the attacker. Weights are assigned to all nodes based on the attacker strategies, with each attacker strategy given equal probability. We develop a novel approximation algorithm for the defender to generate a near-optimal defender strategy. Thus, for the defender, the problem can be considered as finding a near-optimal path with the aim of covering at least one node from each attacker strategy. For this, we use different colors to represent each attacker strategy. The developed defender strategy includes as many different colored vertices as possible. We consider a Stackelberg game instead of a zero-sum game. The advantage of the Stackelberg game over the zero-sum game is that it allows for a more realistic representation of real-world scenarios where one player has more information, resources, or power than the other player. In this Stackelberg game, the defenders have an advantage because they can consider the attacker’s potential moves and responses when making their decisions. This can result in a more efficient outcome than in a zero-sum game, where the players are equally matched.
@@
Due to the complex transportation network and limited police resources, generating efficient strategies for both players is a challenging task. We can summarize the main contributions of this paper as follows: (1) To the best of our knowledge, this is the first attempt to solve a Stackelberg game considering the escape interdiction game (EIG) problem using the multi-layer network concept, with a focus on developing solution methodologies. (2) We develop an optimal approach for the attacker and a near-optimal approach for the defenders, and we are not aware of any previous implementations for escape interdiction problems on multi-layer network. (3) We generate a multi-layer network to create efficient strategies for both players. (4) We use grids dataset for the case study and show that our MLN-EIGS approach outperforms MILP-EIGS approach in terms of solution quality, with a minimal performance gap.
@@
The paper is organized as follows. In Section 2, we present the relevant research. In Section 3, we define the problem description and modeling. In Section 4, we describe the solution methodology. Section 5 presents the quality of the results. We conclude this research in Section 6.
@@ 2 Related Work
Security games play an important role in providing social security. Zhang et al. [3] consider a zero-sum game for the escape interdiction problem and develop a MILP-based solution approach to provide an optimal solution. To demonstrate the efficacy of their methodology, they generate optimal solutions on grids of different sizes. Samanta et al. [1] consider the same zero-sum game problem and develop a meta-heuristic-based solution approach to provide a scalable near-optimal solution in a time-efficient manner. For the same escape interdiction problem, Samanta et al. [2] develop a simulation-based approach to generate a scalable solution to increase security in a large transportation network.
Recent literature focuses on Stackelberg games considering the escape interdiction problem to decrease the crime rate in society. For example, Zychowski and Mandziuk [4] consider search games (SEG) on directed graphs. They consider multiple defender resources and one attacker, where the attacker aims to reach one of several predefined target points from a fixed location. They develop a genetic algorithm-based heuristic approach to provide a near-optimal solution on synthetic datasets.
@@
Basilico et al. [9] consider a leader-follower game and formulate the problem as a mathematical programming model. They use optimization software tools to solve the model by generating the optimal strategies for both leader and follower. Letchford and Conitzer [11] consider security games on graphs and develop a polynomial-time algorithm to generate optimal strategies for players. Similarly, Iwashita et al. [10] consider security games on graphs and develop an algorithm to reduce the graph by eliminating unnecessary edges, providing a time-efficient, scalable near-optimal solution. In the same vein, Karwowski and Mandziuk [18] consider an evader-defender Stackelberg game model and develop a Monte Carlo Tree Search approach to provide efficient patrolling schemes.
@@
Tsai et al. [13] consider an attacker-defender Stackelberg game model and develop a linear program to generate optimal mixed strategies for players by allocating limited resources optimally, with a case study on part of the Mumbai road network. Similarly, Shieh et al. [12] develop a game-theoretic system to provide security with limited resources in the port of Boston and also test its efficacy in the port of New York. They schedule patrols efficiently to provide optimal mixed strategies for players, considering an attacker-defender Stackelberg game model. Similar papers focus on Stackelberg security games to provide security in society (e.g., Cermak et al. [16], Lou et al. [19], Sinha et al. [14], ˇCern‘y et al. [17], Zhang and Malacaria [15]).
@@
Conitzer and Sandholm [5] show that exact approaches are not suitable because of the NP-hard nature of this problem, as it faces scalability issues due to the high time complexity. Paruchuri et al. [6] develop a MILP-based exact algorithm to solve Bayesian Stackelberg security games. Bosansky and Cermak [7] also consider Stackelberg games and develop algorithms based on LPs and MILP to generate Strong Stackelberg Equilibrium (SSE) and perform a case study at Los Angeles International Airport, considering security scheduling.
@@
The layered graph concept is a useful tool for solving complex problems on transportation networks in a time-efficient manner. For example, Saito et al.[8] consider the problem of selecting important nodes in a network and construct a layered graph from the original graph, where each layer is added on top as time proceeds to demystify the complex problem. Hence, we focus on solving a Stackelberg escape interdiction game using the layered graph concept.
@@ 3 Problem Description
We consider a two-player Stackelberg game where this sequential game is played between multiple defenders $(D)$ and one attacker $(A)$. The defenders have a finite set of actions $\acute{S}$, and the attacker has a finite set of actions represented by $\acute{A}$. $S$ and $A$ represent the sets of probability distributions over the defender and attacker actions, respectively. The game is modeled as a general-sum game, where $x\in S$ and $y\in A$ are the probability distributions over $\acute{S}$ and $\acute{A}$, respectively.
For $a_{j}=(v_{j},t^{a}_{j})$ and $s^{r}_{i}=(v^{r}_{i},t^{r,in}_{i},t^{r,out}_{i})$, $d_{r}$ reaches the attacker at node $v^{r}_{i}$ if $v^{r}_{i}$ = $v_{j}$ and $t^{r,in}_{i}\leq t^{a}_{j}\leq t^{r,out}_{i}$. The defender gets -1 if the attacker successfully escapes from the transportation network; otherwise, the defender’s utility value is 0 (see Eq. 1). Considering this case, we define $Z_{s^{r}_{i},a_{j}}=1$, otherwise $Z_{s^{r}_{i},a_{j}}=0$.
@@
The expected utility of the defender and attacker are $U_{d}$ and $U_{a}$, respectively. For any pair of strategies $(x,y)$, the expected payoff of the defender is denoted by $U_{d}(x,y)$.
First, in Eq. (7), the defender commits to a strategy $x\in X$, and given such an $x$, the attacker chooses his strategy from the best-response set $BR(x)$ where
@@
Second, in Eq. (8), to maximize utility, the defender chooses the strategy $x$ (the best response of the defender) to the attacker’s best response.
@@
Third, in a Strong Stackelberg Equilibrium (SSE), we assume that the follower (attacker) breaks ties in favor of the leader (defender) (see Eq. 9). In that case, the optimization problem is
@@
SSE is popular because it is always guaranteed to exist ((Conitzer [20]).
@@ 4 Proposed Mixed Approach on Multi-Layer Networks
This section presents the proposed methodology for the considered escape interdiction problem which is formulated as a Stackelberg game model. A novel mixed approach on layered graph concept is developed to solve the Stackelberg model, where the defender is the leader and the attacker is the follower. For the defender, the developed approximation algorithm finds the efficient movement of the defender from a random initial location. For the attacker, the initial strategy is any random path from the crime location to any exit point. An exact algorithm is developed to generate the optimal path for the attacker using Dijkstra’s algorithm on the time-expanded network. The new defender strategy is generated using a greedy algorithm on the time-expanded network. The defender’s mixed strategy set is updated based on the recalculated utility. The defender then selects the best strategy to commit to. This best defender strategy is given as input to the attacker. The attacker tries to generate a new optimal strategy using the exact approach, taking the new defender strategy as input. The attacker updates the pure attacker strategy set, where all strategies are of equal probability. The algorithm checks if, for $K$ consecutive iterations, the game utility is below a threshold value named $\epsilon$. The value of $\epsilon$ is set to 0.05 for our experiments. If this convergence criterion is met, the algorithm converges, and we return the best defender strategy with the probability of interdiction and total runtime (see Algorithm 1). The flowchart is shown in Fig. 1.
Figure 1: Flowchart of Stackelberg Game Model
@@
Algorithm 1 MLN-EIGS Algorithm
To create a benchmark, we construct a novel Stackelberg game where the defender is the leader, and the attacker is the follower. The best strategies for the attacker and defender are generated using bestAO and bestDo, which are the MILP approaches for the attacker and defender, respectively (see Algorithm 2). Zhang et al. [3] develop these optimal approaches to provide optimal strategies for attackers and defenders given a strategy set for each player. Similar to the MLN-EIGS, in this MILP-EIGS the defender commits to the best mixed strategy. The attacker takes this best defender mixed strategy as input, generates the optimal attacker strategy, and updates the pure attacker strategy set. Then, we calculate the game utility, and the game continues until the game utility is below $\epsilon$ for $K$ consecutive iterations. Here $\epsilon$ acts as a threshold with a value of 0.05. We return the defender utility value once the convergence is reached.
Output: $U_{d}:$ Defenders’ game utility;
repeat
       $(x,y)\leftarrow CoreLP(S^{\prime},A^{\prime})$;
       $BR:\;A^{*}\leftarrow{BestAO(x^{\star})}$;
       if $A^{*}\neq\emptyset$ then
             $A{{}^{\prime}}\leftarrow A^{{}^{\prime}}\cup A^{*}$;
       end if
      $S^{*}\leftarrow{BestDO(A^{{}^{\prime}})}$;
       if $S^{*}\neq\emptyset$ then
             $S^{{}^{\prime}}\leftarrow S^{{}^{\prime}}\cup S^{*}$;
       end if
until convergence;
return $U_{d}$.
Algorithm 2 MILP-EIGS Benchmarking Algorithm
@@ 4.1 Efficient Attacker Strategy Design Using Exact Approach on Time Expanded Network
Algorithm 3 Optimal Attacker Strategy Design Using Time Expanded Network
@@ 4.2 Efficient Defender Strategy Design Using Approximation Algorithm on Time Expanded Network
    Here, each colored vertex corresponds to a particular attacker’s strategy (a path).
@@
    To find a path whose length is at most $t$ and contains as many different colored vertices as possible.
@@
    This problem is difficult as Dijkstra cannot solve its decision version, which is $NP$-complete. We provide a formal proof in this paper.
We can use the steps below for the approximate algorithm.
@@
    Step1: Find a shortest path to any one of colored vertices from the start vertex (we can use Dijkstra). Assume the path is to vertex $v_{1}$ with color $c_{i_{1}}$.
@@
In a simplest form, if the path we obtained in the previous method does not cover a subset of colors, we create another path from the initial vertex, which tries to cover these remaining colors only. We repeat this procedure until all colors are covered. Then, the defender flips a coin and chooses one path.
4.3 Theorem: Problem ”color covering” is $NP$-complete
Problem ”color covering”

### v2

@@ Abstract
Interdicting a criminal with limited police resources is a challenging task in a dynamic crime scenario, as the criminal changes location over time. The vastness of the transportation network adds to the difficulty. To address this, we introduce the concept of a layered graph, where at each time step, a duplicate of the transportation network is generated to trace the possible movements of both the criminal/attacker and the police/defenders. We model this as a Stackelberg game, where the attacker seeks to escape from the network using one of the predefined exit points, while the defenders attempt to intercept the attacker on his escape route. Given a set of defender strategies, the optimal attacker strategy is determined by applying Dijkstra’s algorithm on the layered networks. The attacker seeks to minimize, while the defenders aim to maximize the probability of interdiction. We propose an approximation algorithm on the layered networks to find near-optimal strategy for defenders. The efficacy of the developed approach is compared with the adopted MILP approach. We compare the results in terms of computational time and solution quality. The quality of the results underscores the necessity of the developed approach, as it efficiently solves this complex problem in a short timeframe.
@@
Resource allocation, Multi-layer time expanded network, Vehicle routing and scheduling, Exact algorithm, Approximation algorithm
@@ 1 Introduction
We consider a Stackelberg game where the defender is the leader, and the attacker is the follower. The assumption is that the defenders have information only about the crime location in a large transportation network. The defenders attempt to capture the attacker before he can flee the city. A novel mixed approach on layered graph concept is developed to solve the Stackelberg model. For the defender, the developed approximation algorithm finds the efficient movement of the defender from a random initial location. For the attacker, the initial strategy is any random path from the crime location to any exit point. An exact algorithm is developed to generate the optimal path for the attacker using Dijkstra’s algorithm on the time-expanded network. The new defender strategy is generated using a greedy algorithm on the time-expanded network. The defender’s mixed strategy set is updated based on the recalculated utility. The defender then selects the best strategy to commit to. This best defender strategy is given as input to the attacker. The attacker tries to generate a new optimal strategy using the exact approach, taking the new defender strategy as input. The attacker updates the pure attacker strategy set, where all strategies are of equal probability. The algorithm checks if, for a certain number of consecutive iterations, the game utility is below a threshold value. If this convergence criterion is met, the algorithm converges, and we return the best defender strategy with the probability of interdiction and total runtime. To set a benchmark, the developed Stackelberg game is solved using the adopted MILP approaches for both the attacker and defender.
@@
We consider a Stackelberg game instead of a zero-sum game. The advantage of the Stackelberg game over the zero-sum game is that it allows for a more realistic representation of real-world scenarios where one player has more information, resources, or power than the other player. In this Stackelberg game, the defenders have an advantage because they can consider the attacker’s potential moves and responses when making their decisions. This can result in a more efficient outcome than in a zero-sum game, where the players are equally matched.
@@
Due to the complex transportation network and limited police resources, generating efficient strategies for both players is a challenging task. We can summarize the main contributions of this paper as follows: (1) To the best of our knowledge, this is the first attempt to solve a Stackelberg game considering the escape interdiction game (EIG) problem using the multi-layer network concept, with a focus on developing solution methodologies. (2) We develop an optimal approach for the attacker and a near-optimal approach for the defenders, and we are not aware of any previous implementations for escape interdiction problems on multi-layer network. (3) Using grid datasets, we demonstrate that our method significantly outperforms the MILP approach in both computational time and solution quality, with minimal performance trade-offs.
@@
The paper is organized as follows. In Section 2, we present the relevant research. In Section 3, we define the problem description and modeling. In Section 4, we describe the solution methodology. The benchmarking algorithm is described in Section 5. Section 6 presents the quality of the results. We conclude this research in Section 7.
@@ 2 Related work
Security games play an important role in providing social security ([23], [25]). Recent literature focuses on Stackelberg games considering the escape interdiction problem to decrease the crime rate in society. For example, [4] consider search games (SEG) on directed graphs. They consider multiple defender resources and one attacker, where the attacker aims to reach one of several predefined target points from a fixed location. They develop a genetic algorithm-based heuristic approach to provide a near-optimal solution on synthetic datasets. Similarly, [9] consider a leader-follower game and formulate the problem as a mathematical programming model. They use optimization software tools to solve the model by generating the optimal strategies for both leader and follower. Again, [11] consider security games on graphs and develop a polynomial-time algorithm to generate optimal strategies for players. Similarly, [10] consider security games on graphs and develop an algorithm to reduce the graph by eliminating unnecessary edges, providing a time-efficient, scalable near-optimal solution. In the same vein, [18] consider an evader-defender Stackelberg game model and develop a Monte Carlo Tree Search approach to provide efficient patrolling schemes.
@@
[13] consider an attacker-defender Stackelberg game model and develop a linear program to generate optimal mixed strategies for players by allocating limited resources optimally, with a case study on part of the Mumbai road network. Likewise, [12] develop a game-theoretic system to provide security with limited resources in the port of Boston and also test its efficacy in the port of New York. They schedule patrols efficiently to provide optimal mixed strategies for players, considering an attacker-defender Stackelberg game model. Similar papers focus on Stackelberg security games to provide security in society (e.g., [16], [19], [14], [17], [15]).
@@
[5] show that exact approaches are not suitable due to the NP-hard nature of this problem, as it faces scalability issues due to the high time complexity. On the contrary, [6] develop a MILP-based exact algorithm to solve Bayesian Stackelberg security games. Similarly, [7] consider Stackelberg games and develop algorithms based on LPs and MILP to generate Strong Stackelberg Equilibrium (SSE) and perform a case study at Los Angeles International Airport, focusing on security scheduling.
Considering a zero-sum game for the escape interdiction problem, [3] develop a MILP-based solution approach to provide an optimal solution. To demonstrate the efficacy of their methodology, they generate optimal solutions on grids of different sizes. [1] consider the same zero-sum game problem and develop a meta-heuristic-based solution approach to provide a scalable near-optimal solution in a time-efficient manner. Again, [2] develop a simulation-based approach to generate a scalable solution to increase security in a large transportation network for this escape interdiction problem.
@@
The layered graph concept is a useful tool for solving complex problems on transportation networks in a time-efficient manner. For example, [8] consider the problem of selecting important nodes in a network and construct a layered graph from the original graph, where each layer is added on top as time proceeds to demystify the complex problem. We focus on solving a Stackelberg escape interdiction game using the layered graph concept.
@@ 3 Problem description
We consider a two-player Stackelberg game where this sequential game is played between multiple defenders $(D)=\{d_{r}\mid r\in R\}$ and one attacker $(A)$. The defenders have a finite set of actions $\acute{S}$, and the attacker has a finite set of actions represented by $\acute{A}$. $S$ and $A$ represent the sets of probability distributions over the defender and attacker actions, respectively. The game is modeled as a general-sum game, where $x\in S$ and $y\in A$ are the probability distributions over $\acute{S}$ and $\acute{A}$, respectively. The game begins at time 0 and ends at time $t_{max}>0$. $v_{\infty}$ signifies any exit node in the considered network.
The sequence of states $\textit{A}=<a_{1}=(v^{a}_{0},0),...,a_{j}=(v_{j},t^{a}_{j}),...,a_{k}=(v_{\infty},t^{a}_{k}\leq t_{max})>$ signifies the attacker’s pure strategy. Each state $a_{j}=(v_{j},t^{a}_{j})$ indicates that at time $t^{a}_{j}$, the attacker is present at node $v_{j}$. Likewise, a defender’s state $d_{r}$ is a tuple $S^{r}=(v^{r},t^{r,in},t^{r,out})$, representing the state where defender $d_{r}$ is present at the node $v^{r}$ during the interval $[t^{r,in},t^{r,out}]$.
The defender’s pure strategy consists of a set of $m$ schedules corresponding to the total number of defenders, $m$. The schedule for defender $d_{r}$ is defined as a sequence of states $S^{r}=<s^{r}_{1},...,s^{r}_{i},...,s^{r}_{k}>$, where $s^{r}_{1}=(v^{r}_{0},0,t^{r,out}_{1})$ and $s^{r}_{k}=(v^{r}_{k},t^{r,in}_{k},t_{max})$. The mixed strategy for the defender is denoted by $x=<x_{S}>$, where $x_{S}$ represents the probability with which the strategy $S$ is employed.
For $a_{j}=(v_{j},t^{a}_{j})$ and $s^{r}_{i}=(v^{r}_{i},t^{r,in}_{i},t^{r,out}_{i})$, the defender $d_{r}$ intercepts the attacker at node $v^{r}_{i}$ if $v^{r}_{i}$ = $v_{j}$ and $t^{r,in}_{i}\leq t^{a}_{j}\leq t^{r,out}_{i}$. In this context, we define $z_{s^{r}_{i},a_{j}}=1$ if the conditions are met, otherwise $z_{s^{r}_{i},a_{j}}=0$. The defender incurs a utility of -1 if the attacker successfully escapes from the transportation network; otherwise, the defender’s utility is 0 (see Eq. 1).
@@
The expected utility of the defender and attacker are $U_{d}$ and $U_{a}$, respectively (see Eqs. 7-8). For any pair of strategies $(x,y)$, the expected payoff of the defender is denoted by $U_{d}(x,y)$.
First, considering that the defender commits to a strategy $x\in X$, and given such an $x$, the attacker chooses his strategy from the best-response set $BR(x)$ where
@@
Second, to maximize utility, the defender chooses the strategy $x$ (the best response of the defender) to the attacker’s best response.
@@
Third, in a SSE, we assume that the follower (attacker) breaks ties in favor of the leader (defender) (see Eq. 9). In that case, the optimization problem is
@@
SSE is popular because it is always guaranteed to exist ([21]). The flowchart of the Stackelberg game model is shown in Fig. 1.
Figure 1: Flowchart of Stackelberg game model.
@@ 4 Proposed mixed approach on multi-layer networks
This section presents the proposed methodology for the considered escape interdiction problem which is formulated as a Stackelberg game model. A novel mixed approach on layered graph concept is developed to solve the Stackelberg model, where the defender is the leader and the attacker is the follower. In this Stackelberg game, first, the defenders commit to a strategy, and then the attacker generates the best response to the given defender strategy. We use the concept of a multi-layer network (MLN) in which, at each time-stamp, we create a copy of the entire network. We consider edge length as the time factor to create connections between these multi-layer networks. This means that depending on the edge length, we choose the layers from which the start and end nodes of that particular edge are selected. In this way, for all edges in the original transportation network, we create corresponding edges in the multi-layer network. We consider a set of strategies for both players, i.e., the defenders and the attacker. For the defender, we use a mixed strategy set where each strategy is assigned a mixed probability, with the sum of these probabilities equaling one. For the attacker, we consider a pure strategy set.
To generate the optimal attacker strategy, we first assign weights to the nodes depending on the mixed probabilities and the nodes’ presence in the defender strategies. We check the presence of each node in all defender strategies. If the node is present in multiple strategies, we sum the mixed probabilities of those strategies. This final probability is then assigned as the weight of the node. Then, we assign the weight of each node to all of its incoming edges. Finally, we apply Dijkstra’s algorithm on this network to find the optimal strategy for the attacker. For the defender, we create a multi-layer network in the same way as for the attacker. Weights are assigned to all nodes based on the attacker strategies, with each attacker strategy given equal probability. We develop a novel approximation algorithm for the defender to generate a near-optimal defender strategy. Thus, for the defender, the problem can be considered as finding a near-optimal path with the aim of covering at least one node from each attacker strategy. For this, we use different colors to represent each attacker strategy. The developed defender strategy includes as many different colored vertices as possible. The algorithm converges if for $K$ consecutive iterations, the game utility is below a threshold value named $\epsilon$. Then the best defender strategy with the maximum probability of interdiction and total runtime are returned (see Algorithm 1).
@@
Algorithm 1 MLN-EIGS algorithm.
@@ 4.1 Efficient attacker strategy design using exact approach on time expanded network
Algorithm 2 Optimal attacker strategy design using time expanded network.
@@ 4.2 Efficient defender strategy design using approximation algorithm on time expanded network
    Here, each colored vertex corresponds to a particular attacker’s strategy.
@@
    To find a defender’s strategy whose length is at most $t$ and contains as many different colored vertices as possible.
@@
    This problem is difficult (say, its decision version is $NP$-complete) and thus cannot be solved by Dijkstra. We provide a formal proof in this paper.
We use the steps below for the approximate algorithm.
@@
    Step1: Find a shortest path to any one of colored vertices from the start vertex using Dijkstra. Assume the path is to vertex $v_{1}$ with color $c_{i_{1}}$.
@@
In the simplest form, if the path we obtained in the previous method does not cover a subset of colors, we create another path from the initial vertex, which tries to cover these remaining colors only. We repeat this procedure until all colors are covered. Then, the defender flips a coin and chooses one path.
4.3 Problem ‘color covering’ is $NP$-complete
Problem ‘color covering’